# Stability of vehicular admission control schemes in urban traffic networks under modelling uncertainty

### v1

@@ V-D Second-order linear control dynamics
Depending on the values of the parameters, if ($\mathbf{A}$, $\mathbf{B}$) are controllable, and assuming that (19) holds, the conditions given in 2, and 3 can be efficiently verified by means of a linear matrix inequality (LMI) using the KYP Lemma[KHALIL]. Specifically, if there exist matrices $\mathbf{P}\in\mathbb{R}^{3\times 3},\mathbf{M}\in\mathbb{S}^{2}$ such that
@@
with $\mathbf{Q},\mathbf{R}\in\mathbb{S}^{1},\mathbf{S}\in\mathbb{R}$, then the feedback interconnection is strictly passive with a storage function equal to $V=\boldsymbol{x}_{\rho_{i}}^{T}\mathbf{M}\boldsymbol{x}_{\rho_{i}}$[Scherer2011LinearMI]. As an example, we used CVX[cvx, gb08] to find $\mathbf{P}$, and $\mathbf{M}$ for the parameters documented in Table II (see VAC scheme 4). The solver yielded
@@ V-F VAC dynamics that respect upper and lower bound input constraints
The next scheme to be considered is motivated by [kasis2017primary] and allows to ensure that $u_{i}(t)\in[u_{i}^{\min},u_{i}^{\max}]$. Consider the following input filter
$p^{c}_{i}(\rho_{i})=\begin{cases}p^{c,max}_{i},\;\;\rho_{i}<\rho^{th,1}_{i}\\ -\gamma^{c}_{i}(\rho_{i}-\rho_{i}^{th,1})+p^{c,max}_{i},\;\rho^{th,1}_{i}\leq\rho_{i}<\rho^{th,2}_{i}\\ p^{c,min}_{i},\;\;\rho_{i}\geq\rho^{th,2}_{i}\end{cases}$ (31)
@@
where $\beta^{c}_{i}>0$ and $u_{3,i}$ is a design constant. Moreover, $G_{i}(s)$ describes the dynamics of the system with input $p^{c}_{j}$ and output $u_{1,i}$ in the Laplace domain, as
$G_{i}(s)=K_{i}\frac{(1+sT_{1,i})}{(1+sT_{2,i})(1+sT_{3,i})},$ (33)
where $K_{i},T_{1,i},T_{2,i}$, and $T_{3,i}$ are positive constants. For this system, consider that $\rho_{i}\in(\tilde{\rho}_{i},\hat{\rho}_{i})$, e.g., $\tilde{\rho}_{i}=0$ and $\hat{\rho}_{i}=\rho^{J}_{i}$. These dynamics can be designed such that $u_{i}(t)\in[u_{i}^{\min},u_{i}^{\max}],t\geq 0$. In particular, it can be shown that the trajectories of $u_{i}(t)$ can be bounded by evaluating the $\mathcal{L}_{1}$ norm of $G_{i}$ and noting that both $\rho_{i}$ and $p^{c}_{i}$ lie in bounded sets. Then, given the description of $u_{1,i}$ and $u_{2,i}$, trajectory bounds may be obtained through suitable design of the constants $p^{c,min}_{i},p^{c,max}_{i},\gamma^{c}_{i},$ and $u_{3,i}$. Conditions that enable the passivity of the above system are presented in the following lemma.
@@
Next, the VAC dynamics examples from Section V will be employed on a representative simulation of a 6-region vehicular traffic network to demonstrate the advantages of the proposed approach.
@@ VI Simulation Results
This section presents numerical simulations, conducted on a 6-region network and a 20-region network, under varying disturbance and uncertainty conditions. The main result of this work, 1, is demonstrated through both networks, while the scalability and robustness of the proposed approach under uncertainty are demonstrated through simulations on the 20-region network. The section concludes with a brief discussion on simulations that explore the sensitivity of the analytic stability result to controller parameter variations.
@@
TABLE III: 6-region network parameters[MENELAOU2023]
@@ VI-C Demonstrating Scalability
The following numerical simulation aims to demonstrate the scalability and large scale applicability of the developed approach. Specifically, to demonstrate that the presented results are also applicable to larger networks, the simulation from Section VI-A was repeated on the 20-region network shown in Fig. 12(a). In this simulation experiment, (21), (23), (24), and (26) are employed for regions 1 to 5, 6 to 10, 11 to 15, and 16 to 20, respectively.
@@
TABLE VI: 20-region network parameters[MENELAOU2023]
@@
(21) $\boldsymbol{\eta}_{1,..,5}=[284.3,277.1,351.1,288.1,206.2]^{T}$
$\mathbf{c}_{1,..,5}=[6783.1,4017.3,2300.4,2544.2,2070.1]^{T}$
$\boldsymbol{\upsilon}_{1,..,5}=[0.001,0.001,0.001,0.001,0.001]^{T}$
(23) $\boldsymbol{\eta}_{6,..,10}=[344.4,318.6,332.9,252,325.6]^{T}$
$\mathbf{c}_{6,..,10}=[5672.5,8468.4,4031.0,6472.9,3816.3]^{T}$
$\phi_{i}(\rho_{i})=0.001\rho_{i}^{3},i=6,..,10$
(24) $\boldsymbol{\eta}_{11,..,15}=[228.4,199.2,172.3,214.7,176.5]^{T}$
$\boldsymbol{\gamma}_{11,..,15}=[114.2,99.6,86.1,107.4,88.3]$
$\mathbf{c}_{11,..,15}=[3909.4,6541.1,4224.1,3310.7,9054.5]^{T}$
$\tau_{i}=0.1,i=11,..,15$
(26) $\boldsymbol{\eta}_{16,..,20}=[314.3,265.6,231.1,230.1,214.1]^{T}$
$\mathbf{c}_{16,..,20}=[2146.2,5283.6,4775.0,6345.3,5776.1]^{T}$
$\tau_{i}=0.1,\kappa_{i}=0.2,i=16,..,20$
@@
(21) $\boldsymbol{\eta}_{1,..,5}=[284.3,277.1,351.1,288.1,206.2]^{T}$
$\mathbf{c}_{1,..,5}=[5426.5,3213.6,1840.3,2035.4,1656.1]^{T}$
$\boldsymbol{\upsilon}_{1,..,5}=[0.001,0.001,0.001,0.001,0.001]^{T}$
(23) $\boldsymbol{\eta}_{6,..,10}=[344.4,318.6,332.9,252,325.6]^{T}$
$\mathbf{c}_{6,..,10}=[4536.7,6768.9,3224.3,5173.7,3052.6]^{T}$
$\phi_{i}(\rho_{i})=0.001\rho_{i}^{3},i=6,..,10$
(24) $\boldsymbol{\eta}_{11,..,15}=[228.4,199.2,172.3,214.7,176.5]^{T}$
$\boldsymbol{\gamma}_{11,..,15}=[114.2,99.6,86.1,107.4,88.3]$
$\mathbf{c}_{11,..,15}=[3126.9,5236.0,3378.8,2648.3,7243.1]^{T}$
$\tau_{i}=0.1,i=11,..,15$
(26) $\boldsymbol{\eta}_{16,..,20}=[314.3,265.6,231.1,230.1,214.1]^{T}$
$\mathbf{c}_{16,..,20}=[1717,4226,3819.7,5076.5,4621.5]^{T}$
$\tau_{i}=0.1,\kappa_{i}=0.2,i=16,..,20$
@@ VI-D Assessing Network Stability Under Uncertainty
Next, to demonstrate the robustness of the developed VAC framework while operating under uncertainty, 200 different simulations are conducted on the 20-region network from Section VI-C as follows. The simulation scenario from Section VI-A is repeated 200 times, each time with a different randomly generated uncertainty function, $d_{i}(\rho_{i})$, abiding by the same conditions, as presented in 1. The same fixed control gains are used across all simulations (shown in Table VIII), allowing to demonstrate the framework’s ability to handle a range of perturbation functions with constant gain parameters. In analogy with Section VI-C, the VAC schemes (21), (23), (24), and (26) are employed for regions 1 to 5, 6 to 10, 11 to 15, and 16 to 20, respectively. The procedure for assessing network stability under uncertainty is outlined below.
@@
The series of experiments regarding robustness are presented next. As depicted in Fig. 16, fast steady-state network stabilization is observed in the first 30 seconds. At $t=30\,\text{[min]}$, the developed VAC framework becomes disengaged (for $1.5\,\text{[min]}$) and unregulated vehicles enter the network from regions 1, 3, and 5, at capacity flow rates, making some regions congested (regions 7, 9, 12 and 15, see Fig. 14(b)) and resulting in increased error.
When the VAC framework is re-engaged it reduces vehicle inflow to congested regions and increases vehicle inflow to less dense regions. As a result, upon re-engagement, the error to non-congested regions quickly decreases. In contrast, due to the increased inflow of vehicles to non-congested regions, inter-regional flows to connecting congested regions also increase, momentarily worsening congestion levels despite the reduction of vehicle inflow from outside the network. Thus the resulting error momentarily increases after it first drops. Despite this behaviour, in all the conducted simulations the presented results, as demonstrated in Fig. 16, validate the ability of the developed VAC design framework to stabilize the system, even at the presence of significant uncertainty and control disruptions.
@@ VI-E Assessing VAC scheme sensitivity to parameter changes
To evaluate the ability of different VAC dynamic schemes to regulate demand in a stable manner, as well as their sensitivity to changes in VAC parameters, we conducted 50 simulations on the 20-region network from Section VI-A using VAC dynamics with randomly generated parameters that satisfy the presented stability conditions. The desired operating point for this series of simulations is given in Table IV. Similar to the previous sections, the VAC schemes (21), (23), (24), and (26) are employed for regions 1 to 5, 6 to 10, 11 to 15, and 16 to 20, respectively. An indicative set of random gain values employed in one of the simulations, which satisfies the conditions of 1, is reported in Table IX. To condense the results into a single plot, the same method employed in Section VI-D is used to calculate the worst maximum density deviation from the desired operating point. Pointedly, for each simulation the absolute density deviation from the desired operating point, $e_{\rho_{i}}^{*}(t)=|\rho_{i}(t)-\rho_{i}^{*}|,i\in\mathcal{N}$, is calculated for each region. Using $e_{\rho_{i}}^{*}$, the maximum density deviation across all regions at each time step is then recorded as $e_{\rho}^{*}(t)=\max_{i\in\mathcal{N}}\{e_{\rho_{i}}^{*}(t)\}$. Finally, for each time interval, the worst maximum density deviation across the 50 simulations, $e_{\rho,50}^{*}(t)=\max_{1,..,50}\{e_{\rho}^{*}(t)\}$, is saved into a new time series.
@@
(21) $\boldsymbol{\eta}_{1,..,5}=[505.7,338.6,445.6,271.9,422.2]^{T}$
$\mathbf{c}_{1,..,5}=[6570,4731,4256,5461,10452]^{T}$
$\boldsymbol{\upsilon}_{1,..,5}=[0.0017,0.0016,0.0017,0.0015,0.0017]^{T}$
(23) $\boldsymbol{\eta}_{6,..,10}=[548.4,321.4,563.9,607.3,538.3]^{T}$
$\mathbf{c}_{6,..,10}=[11336,11791,13569,9825,18960]^{T}$
$\boldsymbol{\varrho}_{6,..,10}=[0.0013,0.0014,0.0018,0.0016,0.0016]^{T}$
$\phi_{i}(\rho_{i})=\varrho_{i}\rho_{i}^{3},i=6,..,10$
(24) $\boldsymbol{\eta}_{11,..,15}=[247.7,259.9,406.5,304,218.2]^{T}$
$\boldsymbol{\gamma}_{11,..,15}=[123.9,129.9,203.3,152,109.1]$
$\mathbf{c}_{11,..,15}=[9360,6057,7682,9885,9264]^{T}$
$\boldsymbol{\tau}_{11,..,15}=[0.0925,0.0892,0.0785,0.0577,0.0285]^{T}$
(26) $\boldsymbol{\eta}_{16,..,20}=[512.1,625.6,288.6,378.3,349.1]^{T}$
$\mathbf{c}_{16,..,20}=[10162,16442,7501,8023,8248]^{T}$
$\boldsymbol{\tau}_{16,..,20}=[0.0524,0.0509,0.0958,0.0802,0.0378]^{T}$
$\boldsymbol{\kappa}_{16,..,20}=[0.0957,0.1938,0.0504,0.1512,0.1034]^{T}$
The results are presented next. At the initial stages of the simulations (up to $t=30\,\text{[min]}$), fast stabilization to the desired operating point is demonstrated in all cases, see Fig. 17. The varying VAC parameters result in different transient responses for each simulation. At $t=30\,\text{[min]}$ due to the control disruption event the VAC framework is disengaged from all the regions for $1.5\,\text{[min]}$, and unregulated vehicles enter the network from regions 1, 3 and 5 at capacity flow rates causing congestion to some regions. Upon re-engagement of the VAC framework, fast recovery to the desired operating point is observed. The error to non-congested regions quickly decreases. Due to the increased inflow of vehicles to non-congested regions that result to increased inter-regional flow of vehicles to connecting congested regions, the error momentarily increases worsening congestion levels despite the minimization of vehicle inflow from outside the network due to the VAC frameworks employed at congested regions. Thus resulting to the error increasing after it first drops. Despite this behaviour, in all the conducted simulations the presented results, as demonstrated in Fig. 17, validate the ability of the developed VAC design framework to stabilize the system, even at the presence of significant uncertainty and control disruptions. All simulations produced stable outcomes as expected. These results indicate that the proposed VAC framework can effectively regulate vehicular flow, provided that the stability criterion is satisfied.

### v2

@@ V-D Second-order linear control dynamics
Depending on the values of the parameters, if ($\mathbf{A}$, $\mathbf{B}$) are controllable, and assuming that (19) holds, the conditions given in 2, and 3 can be efficiently verified by means of a linear matrix inequality (LMI) using the KYP Lemma[25]. Specifically, if there exist matrices $\mathbf{P}\in\mathbb{R}^{3\times 3},\mathbf{M}\in\mathbb{S}^{2}$ such that
@@
with $\mathbf{Q},\mathbf{R}\in\mathbb{S}^{1},\mathbf{S}\in\mathbb{R}$, then the feedback interconnection is strictly passive with a storage function equal to $V=\boldsymbol{x}_{\rho_{i}}^{T}\mathbf{M}\boldsymbol{x}_{\rho_{i}}$[45]. As an example, we used CVX[14, 13] to find $\mathbf{P}$, and $\mathbf{M}$ for the parameters documented in Table II (see VAC scheme 4). The solver yielded
@@ V-F VAC dynamics that respect upper and lower bound input constraints
The next scheme to be considered is motivated by [21] and allows to ensure that $u_{i}(t)\in[u_{i}^{\min},u_{i}^{\max}]$. Consider the following input filter
$p^{c}_{i}(\rho_{i})=\begin{cases}p^{c,max}_{i},\;\;\rho_{i}<\rho^{th,1}_{i}\\ -\gamma_{i}(\rho_{i}-\rho_{i}^{th,1})+p^{c,max}_{i},\;\rho^{th,1}_{i}\leq\rho_{i}<\rho^{th,2}_{i}\\ p^{c,min}_{i},\;\;\rho_{i}\geq\rho^{th,2}_{i}\end{cases}$ (31)
@@
where $\beta^{c}_{i}>0$ and $c_{i}$ is a design constant. Moreover, $G_{i}(s)$ describes the dynamics of the system with input $p^{c}_{i}$ and output $u_{1,i}$ in the Laplace domain, as
$G_{i}(s)=K_{i}\frac{(1+sT_{1,i})}{(1+sT_{2,i})(1+sT_{3,i})}$ (33)
where $K_{i},T_{1,i},T_{2,i}$, and $T_{3,i}$ are positive constants. For this system, consider that $\rho_{i}\in(\tilde{\rho}_{i},\tilde{\tilde{\rho}}_{i})$, e.g., $\tilde{\rho}_{i}=0$ and $\tilde{\tilde{\rho}}_{i}=\rho^{J}_{i}$. These dynamics can be designed such that $u_{i}(t)\in[u_{i}^{\min},u_{i}^{\max}],t\geq 0$. In particular, it can be shown that the trajectories of $u_{i}(t)$ can be bounded by evaluating the $\mathcal{L}_{1}$ norm of $G_{i}$ and noting that both $\rho_{i}$ and $p^{c}_{i}$ lie in bounded sets. Then, given the description of $u_{1,i}$ and $u_{2,i}$, trajectory bounds, given feasible initial conditions, may be obtained through suitable design of the constants $p^{c,min}_{i},p^{c,max}_{i},\gamma^{c}_{i},$ and $c_{i}$. Conditions that enable the passivity of the above system are presented in the following lemma, proven in the appendix.
@@
Next, the VAC dynamics examples from Section V will be employed on a representative simulation of a 6 and a 20-region vehicular traffic network to demonstrate the advantages of the proposed approach.
@@ VI Simulation Results
This section presents numerical simulations, conducted on a 6-region network and a 20-region network, under varying disturbance and uncertainty conditions. The main result of this work, 1, is demonstrated through the 6-region network, while the scalability and robustness of the proposed approach under uncertainty are demonstrated through simulations on the 20-region network. The section concludes with a brief discussion on simulations that explore the sensitivity of the analytic stability result to controller parameter variations.
@@
TABLE III: 6-region network parameters[36]
@@ VI-C Demonstrating Scalability
The following numerical simulation aims to demonstrate the scalability and large scale applicability of the developed approach. Specifically, to demonstrate that the presented results are also applicable to larger networks, the simulation from Section VI-A was repeated on the 20-region network shown in Fig. 12(a). In this simulation, (21), (23), (24), (26), and (32) are employed for regions 1 to 4, 5 to 8, 9 to 12, 13 to 16 and 17 to 20, respectively.
@@
TABLE VI: 20-region network parameters[36]
@@
(21) $\boldsymbol{\eta}_{1,..,4}=[544.6,454.9,488,428.1]^{T}$
$\mathbf{c}_{1,..,4}=[11775,6411,3086,3660]^{T}$
(23) $\boldsymbol{\eta}_{5,..,8}=[330.8,533.7,613.7,638.7]^{T}$
$\mathbf{c}_{5,..,8}=[3153,8530,15888,7488]^{T}$
$\phi_{i}(\rho_{i})=0.001\rho_{i}^{3},i=5,..,8$
(24) $\boldsymbol{\eta}_{9,..,12}=[234.3,426.2,268.1,224.4]^{T}$
$\boldsymbol{\gamma}_{9,..,12}=[117.2,213.1,134.1,112.2]$
$\mathbf{c}_{9,..,12}=[8794.8,7211.8,4432,7213.1]^{T}$
$\tau_{i}=0.1,i=9,..,12$
(26) $\boldsymbol{\eta}_{13,..,16}=[293.3,485,270.5,612.2]^{T}$
$\mathbf{c}_{13,..,16}=[4702,4834.1,9062.6,3972]^{T}$
$\tau_{i}=0.1,\kappa_{i}=0.2,i=13,..,16$
(32) $\boldsymbol{\rho}^{th,1}_{17,..,20}=[0.4,0.4,0.3,0.3]^{T}$
$\boldsymbol{\rho}^{th,2}_{17,..,20}=[131.4,126.5,118.5,100.8]^{T}$
$\mathbf{p}^{c,max}_{17,..,20}=[134.1,129.1,120.9,102.9]^{T}$
$\boldsymbol{\beta}^{c}_{17,..,20}=[485.3,234.6,270.6,285.2]^{T}$
$\mathbf{c}_{17,..,20}=[9280,4716.1,7259.5,7419.1]^{T}$
$T_{1,i}=0.01,T_{2,i}=0.1,T_{3,i}=0.2,\kappa_{i}=0.1,\gamma_{i}=1.023,i=17,..,20$
$\boldsymbol{\upsilon}_{17,..,20}=[0.001,0.001,0.001,0.001]^{T}$
@@
(21) $\boldsymbol{\eta}_{1,..,4}=[356.8,532.1,457.3,532.6]^{T}$
$\mathbf{c}_{1,..,4}=[6591.9,6111.0,2367.2,3697.7]^{T}$
(23) $\boldsymbol{\eta}_{5,..,8}=[253.6,541.4,378.8,472.7]^{T}$
$\mathbf{c}_{5,..,8}=[2005.4,7059.5,8035.0,4544.2]^{T}$
$\phi_{i}(\rho_{i})=0.001\rho_{i}^{3},i=5,..,8$
(24) $\boldsymbol{\eta}_{9,..,12}=[231.2,225.5,257.6,349.3]^{T}$
$\boldsymbol{\gamma}_{9,..,12}=[115.6,112.8,128.8,174.7]$
$\mathbf{c}_{9,..,12}=[7084.4,3166.8,3488.1,9127.9]^{T}$
$\tau_{i}=0.1,i=9,..,12$
(26) $\boldsymbol{\eta}_{13,..,16}=[365,581.1,442.2,434.6]^{T}$
$\mathbf{c}_{13,..,16}=[4756,4708,11367,2333]^{T}$
$\tau_{i}=0.1,\kappa_{i}=0.2,i=13,..,16$
(32) $\boldsymbol{\rho}^{th,1}_{17,..,20}=[0.4,0.4,0.3,0.3]^{T}$
$\boldsymbol{\rho}^{th,2}_{17,..,20}=[131.4,126.5,118.5,100.8]^{T}$
$\mathbf{p}^{c,max}_{17,..,20}=[134.1,129.1,120.9,102.9]^{T}$
$\boldsymbol{\beta}^{c}_{17,..,20}=[397,450.8,448.2,317.9]^{T}$
$\mathbf{c}_{17,..,20}=[6217.5,7272.2,9773.8,6727.6]^{T}$
$T_{1,i}=0.01,T_{2,i}=0.1,T_{3,i}=0.2,\kappa_{i}=0.1,\gamma_{i}=1.023,i=17,..,20$
$\boldsymbol{\upsilon}_{17,..,20}=[0.001,0.001,0.001,0.001]^{T}$
@@ VI-D Assessing Network Stability Under Uncertainty
Next, to demonstrate the robustness of the developed VAC framework while operating under uncertainty, 200 different simulations are conducted on the 20-region network from Section VI-C as follows. The simulation scenario from Section VI-A is repeated 200 times, each time with a different randomly generated uncertainty function, $d_{i}(\rho_{i})$, abiding by the same conditions, as presented in 1. The same fixed control gains are used across all simulations (shown in Table VIII), allowing to demonstrate the framework’s ability to handle a range of perturbation functions with constant gain parameters. In analogy with Section VI-C, the VAC schemes (21), (23), (24), (26), and (32) are employed for regions 1 to 4, 5 to 8, 9 to 12, 13 to 16 and 17 to 20, respectively. The procedure for assessing network stability under uncertainty is outlined below.
@@
The series of experiments regarding robustness are presented next. As depicted in Fig. 16, fast steady-state network stabilization is observed in the first 30 seconds. At $t=30\,\text{[min]}$, the developed VAC framework becomes disengaged (for $1.5\,\text{[min]}$) and unregulated vehicles enter the network from regions 1, 3, and 5, at capacity flow rates, making some regions congested (regions 7, 9, 12 and 15, see Fig. 14(b)) and resulting in increased error. When the VAC framework is re-engaged it reduces vehicle inflow to congested regions and increases vehicle inflow to less dense regions. As a result, upon re-engagement the error quickly decreases in all the conducted simulations, as demonstrated in Fig. 16, validating the ability of the developed VAC design framework to stabilize the system, even at the presence of significant uncertainty and control disruptions.
@@ VI-E Assessing VAC scheme sensitivity to parameter changes
To evaluate the ability of different VAC dynamic schemes to regulate demand in a stable manner, as well as their sensitivity to changes in VAC parameters, we conducted 50 simulations on the 20-region network from Section VI-A using VAC dynamics with randomly generated parameters that satisfy the presented stability conditions. The desired operating point for this series of simulations is given in Table IV. Similar to the previous sections, the VAC schemes (21), (23), (24), (26), and (32) are employed for regions 1 to 4, 5 to 8, 9 to 12, 13 to 16 and 17 to 20, respectively. An indicative set of random gain values employed in one of the simulations, which satisfies the conditions of 1, is reported in Table IX. To condense the results into a single plot, the same method employed in Section VI-D is used to calculate the worst maximum density deviation from the desired operating point. Pointedly, for each simulation the absolute density deviation from the desired operating point, $e_{\rho_{i}}^{*}(t)=|\rho_{i}(t)-\rho_{i}^{*}|,i\in\mathcal{N}$, is calculated for each region. Using $e_{\rho_{i}}^{*}$, the maximum density deviation across all regions at each time step is then recorded as $e_{\rho}^{*}(t)=\max_{i\in\mathcal{N}}\{e_{\rho_{i}}^{*}(t)\}$. Finally, for each time interval, the worst maximum density deviation across the 50 simulations, $e_{\rho,50}^{*}(t)=\max_{1,..,50}\{e_{\rho}^{*}(t)\}$, is saved into a new time series.
@@
(21) $\boldsymbol{\eta}_{1,..,4}=[717.8,857.9,466.5,429.2]^{T}$
$\mathbf{c}_{1,..,4}=[15187,12019,2955,3669]^{T}$
(23) $\boldsymbol{\eta}_{5,..,8}=[451.3,659.4,1040,550.9]^{T}$
$\mathbf{c}_{5,..,8}=[4240,10501,26829,6472]^{T}$
$\phi_{i}(\rho_{i})=0.001\rho_{i}^{3},i=5,..,8$
(24) $\boldsymbol{\eta}_{9,..,12}=[543.4,523,591.4,313.1]^{T}$
$\boldsymbol{\gamma}_{9,..,12}=[271.7,261.5,295.7,156.6]$
$\mathbf{c}_{9,..,12}=[20289,8820,9327,10032]^{T}$
$\tau_{i}=0.1,i=9,..,12$
(26) $\boldsymbol{\eta}_{13,..,16}=[420.6,703,517.2,475.3]^{T}$
$\mathbf{c}_{13,..,16}=[6697,6949,16049,3113]^{T}$
$\tau_{i}=0.1,\kappa_{i}=0.2,i=13,..,16$
(32) $\boldsymbol{\rho}^{th,1}_{17,..,20}=[0.4,0.4,0.3,0.3]^{T}$
$\boldsymbol{\rho}^{th,2}_{17,..,20}=[131.4,126.5,118.5,100.8]^{T}$
$\mathbf{p}^{c,max}_{17,..,20}=[134.1,129.1,120.9,102.9]^{T}$
$\boldsymbol{\beta}^{c}_{17,..,20}=[535.7,564.3,535.2,550.6]^{T}$
$\mathbf{c}_{17,..,20}=[10228,11114,14287,14103]^{T}$
$T_{1,i}=0.01,T_{2,i}=0.1,T_{3,i}=0.2,\kappa_{i}=0.1,\gamma_{i}=1.023,i=17,..,20$
$\boldsymbol{\upsilon}_{17,..,20}=[0.001,0.001,0.001,0.001]^{T}$
The results are presented next. At the initial stages of the simulations (up to $t=30\,\text{[min]}$), fast stabilization to the desired operating point is demonstrated in all cases, see Fig. 17. The varying VAC parameters result in different transient responses for each simulation. At $t=30\,\text{[min]}$ due to the control disruption event the VAC framework is disengaged from all the regions for $1.5\,\text{[min]}$, and unregulated vehicles enter the network from regions 1, 3 and 5 at capacity flow rates causing congestion to some regions. Upon re-engagement of the VAC framework, fast recovery to the desired operating point is observed, as demonstrated in Fig. 17, validating the ability of the developed VAC design framework to stabilize the system, even at the presence of significant uncertainty and control disruptions. These results indicate that the proposed VAC framework can effectively regulate vehicular flow, provided that the stability criterion is satisfied.